# Counterfactual Visual Explanation via Causally-Guided Adversarial Steering

### v1

@@ B. Additional Experimental Results
Results on CelebA-HQ Due to the similarity between the CelebA-HQ and CelebA datasets, we omit the CelebA-HQ results from the main text and provide them in this section as supplementary material. We first present the quantitative analysis. Tables 7 and 8 summarize the results on the CelebA-HQ dataset, which exhibit a similar overall trend to that observed on CelebA. Although DiME achieves slightly better FR, CD, and sFID scores, it does so at a significant expense of FID. Moreover, DiME is obviously less competitive in terms of perceptual similarity (LPIPS) and cosine similarity (S3). Notably, its COUT score is negative, indicating that the classifier lacks confidence in the counterfactual transitions produced by the DiME.
Table 7: Evaluation Results on CelebA-HQ: Smile.
DiVE - - - - - 107.0 -
STEEX - - - - - 21.9 -
DiME 90.00% -0.2797 0.1689 0.9245 2.52 15.8 26.79
ACE 77.50% 0.2369 0.0353 0.9868 2.76 5.9 38.20
Ours 89.35% 0.3918 0.0190 0.9892 2.74 4.4 27.85
Table 8: Evaluation Results on CelebA-HQ: Age
DiVE - - - - - 107.5 -
STEEX - - - - - 26.8 -
DiME 94.55% -0.3339 0.1689 0.9202 4.00 17.8 26.54
ACE 63.45% 0.0567 0.0448 0.9876 5.73 9.7 33.95
Ours 89.50% 0.3683 0.0278 0.9897 5.53 5.6 27.92
@@
We then present the qualitative analysis. Since ACE achieves the strongest performance among all baselines, we focus our comparison on the counterfactual images generated by our method and those produced by ACE. As shown in Figure 7, similar to the results on the CelebA dataset, the counterfactual images generated by ACE contain many unintended modifications. These include changes in teeth and chin color, the appearance of exaggerated facial creases, and the emergence of extraneous patches on the cheeks. In the transition from “old” to “young,” the counterfactual images generated by our method exhibit comparable visual quality to those produced by ACE, with no obvious undesired changes in either case. As a result, we do not highlight any specific regions for this comparison. However, our method aligns more closely with the intended transformation—for example, by effectively removing age spots and smoothing wrinkles—demonstrating greater semantic consistency with the target concept.
More Results on Other Datasets The extended qualitative results on CelebA, BDD-OIA, and ImageNet are presented in Figures 8,  9, and  10, respectively.
@@
Figure 7: Qualitative results on CelebA-HQ dataset. Red circles highlight unintended changes produced by ACE, while green circles indicate the corresponding improvements made by our method.
@@
Figure 8: Extended qualitative results on CelebA dataset. Red circles highlight unintended changes produced by ACE, while green circles indicate the corresponding improvements made by our method.
@@
Figure 9: Extended qualitative results on BDD-OIA dataset. Green circles highlight the counterfactual changes. The right column for each transition direction shows a zoomed-in version of the left column.
@@
Figure 10: Extended qualitative results on ImageNet dataset. Green circles highlight the counterfactual changes.
C. VLM Prompt Design
@@
CelebA Smile You are given two sets of facial images: one set contains the original images, and the other contains the corresponding counterfactual images. The target attribute to be modified is smile. If the person in the original image is smiling, the target is "not smiling"; if the person is not smiling, the target is "smiling". Please evaluate the counterfactual images based on the following four criteria, each rated on a scale from 1 to 10:
@@
CelebA Age You are given two sets of facial images: one set contains the original images, and the other contains the corresponding counterfactual images. The target attribute to be modified is age. If the person in the original image is old, the target is "young"; if the person is young, the target is "old". Please evaluate the counterfactual images based on the following four criteria, each rated on a scale from 1 to 10:
@@
CelebA-HQ Smile You are given two sets of facial images: one set contains the original images, and the other contains the corresponding counterfactual images. The target attribute to be modified is smile. If the person in the original image is smiling, the target is "not smiling"; if the person is not smiling, the target is "smiling". Please evaluate the counterfactual images based on the following four criteria, each rated on a scale from 1 to 10:
@@
CelebA-HQ Age You are given two sets of facial images: one set contains the original images, and the other contains the corresponding counterfactual images. The target attribute to be modified is age. If the person in the original image is old, the target is "young"; if the person is young, the target is "old". Please evaluate the counterfactual images based on the following four criteria, each rated on a scale from 1 to 10:
@@
Sorrel - Zebra You are given two sets of animal images: one set contains the original images, and the other contains the corresponding counterfactual images. The target attribute to be modified is animal type. The query attribute is "sorrel", and the target attribute is "zebra". Please evaluate the counterfactual images based on the following four criteria, each rated on a scale from 1 to 10:
@@
Cougar - Cheetah You are given two sets of animal images: one set contains the original images, and the other contains the corresponding counterfactual images. The target attribute to be modified is animal type. The query attribute is "cougar", and the target attribute is "cheetah". Please evaluate the counterfactual images based on the following four criteria, each rated on a scale from 1 to 10:
@@
Persian Cat - Egyptian Cat You are given two sets of animal images: one set contains the original images, and the other contains the corresponding counterfactual images. The target attribute to be modified is animal type. The query attribute is "Persian cat", and the target attribute is "Egyptian cat". Please evaluate the counterfactual images based on the following four criteria, each rated on a scale from 1 to 10:
@@
D. Additional Parameters Study
Except for the results on BDD-OIA, the patterns observed in the other datasets are consistent with those presented in the main text. For the BDD-OIA dataset, our method is able to achieve higher LPIPS and COUT scores while maintaining the FR without degradation when $\alpha=0.005$. Although the FR does not decrease within the range of $\alpha$ values considered in Figure 14, it is foreseeable that further increasing $\alpha$ would gradually lead to a decline in FR. On the BDD-OIA dataset, our method is unlikely to achieve a higher FR than ACE. Therefore, further increasing the value of $\alpha$ in pursuit of an ideal trade-off is no longer feasible. The setting $\alpha=0.005$ already yields satisfactory results, which is why we choose it as the final hyperparameter value. We focus on this parameter study—also serving as an ablation when $\alpha=0$—to examine the effect of $\alpha$, while keeping all other hyperparameters consistent with those used in ACE. This ensures a fair comparison and highlights the critical role of $\alpha$ in our method, as it controls the degree of causal constraint on the change of spurious factors.
Figure 11: Line plots of FR and FID with respect to varying values of $\alpha$ on CelebA Age.
Figure 12: Line plots of FR and FID with respect to varying values of $\alpha$ on CelebA-HQ Smile.
Figure 13: Line plots of FR and FID with respect to varying values of $\alpha$ on CelebA-HQ Age.
Figure 14: Line plots of FR and FID with respect to varying values of $\alpha$ on BDD-OIA.
Figure 15: Line plots of FR and FID with respect to varying values of $\alpha$ on ImageNet (Sorrel - Zebra).
Figure 16: Line plots of FR and FID with respect to varying values of $\alpha$ on ImageNet (Cougar - Cheetah.
Figure 17: Line plots of FR and FID with respect to varying values of $\alpha$ on ImageNet (Persian Cat - Egyptian Cat.

### v2

@@ Appendix B Additional Experimental Results
Due to the page limit, we omit the results of CelebA, CelebA-HQ Gender, and BDD-OIA from the main text and provide them in this section as supplementary material. We first present the quantitative analysis. Table 6, Table 7, and Table 8 summarize the results on the CelebA dataset, Table 9 concludes the results on CelebA-HQ Gender, which exhibit a similar overall trend to that observed on CelebA-HQ Smile and Age. Although ACE achieves slightly better performance than our method on LPIPS and S3, this comes at the cost of reduced flip rate and FID. Importantly, FID is a key metric, as it measures whether the overall distribution of counterfactual images is close to that of real images. Similarly, although DiME attains slightly better results on sFID, its COUT is extremely low and even negative, indicating that the classifier lacks confidence in the counterfactual transitions produced by DiME. Table 10 presents the results on the BDD-OIA dataset. Although ACE performs well on several metrics, our method outperforms it in terms of COUT, demonstrating that the classifier retains sufficient confidence in the counterfactual transitions.
Table 6: Evaluation Results on CelebA: Smile.
Methods FR $\uparrow$ COUT $\uparrow$ LPIPS $\downarrow$ S3 $\uparrow$ CD $\downarrow$ FID $\downarrow$ sFID $\downarrow$
DiVE - - - - - 29.4 -
DiVE100 - - - - 2.34 36.8 -
STEEX - - - - - 10.2 -
DiME 97.60% -0.0451 0.0896 0.9158 1.99 12.3 25.48
ACE 99.05% 0.6547 0.0273 0.9822 2.05 4.2 26.38
Ours 99.65% 0.7943 0.0299 0.9819 2.29 3.9 25.70
Table 7: Evaluation Results on CelebA: Age
Methods FR $\uparrow$ COUT $\uparrow$ LPIPS $\downarrow$ S3 $\uparrow$ CD $\downarrow$ FID $\downarrow$ sFID $\downarrow$
DiVE - - - - - 33.8 -
DiVE100 - - - - - 39.9 -
STEEX - - - - - 11.8 -
DiME 73.10% 0.0381 0.0938 0.9158 3.81 20.2 34.37
ACE 99.00% 0.5441 0.0326 0.9843 3.37 4.3 26.47
Ours 99.75% 0.6760 0.0333 0.9829 3.17 4.3 25.76
Table 8: Evaluation results on CelebA: Gender.
Methods FR $\uparrow$ COUT $\uparrow$ LPIPS $\downarrow$ S3 $\uparrow$ CD $\downarrow$ FID $\downarrow$ sFID $\downarrow$
DiME 97.05% 0.0636 0.0914 0.9158 4.91 14.3 27.9
ACE 91.75% 0.4947 0.0360 0.9826 5.44 7.3 29.6
Ours 97.15% 0.6304 0.0387 0.9807 4.53 6.3 27.9
Table 9: Evaluation results on CelebA-HQ: Gender.
Methods FR $\uparrow$ COUT $\uparrow$ LPIPS $\downarrow$ S3 $\uparrow$ CD $\downarrow$ FID $\downarrow$ sFID $\downarrow$
DiME 98.45% -0.3630 0.1671 0.9146 7.55 17.3 28.0
ACE 26.75% -0.3705 0.0376 0.9865 9.66 26.7 55.4
Ours 99.95% 0.8555 0.0627 0.9831 8.30 7.6 29.0
Table 10: Evaluation results on BDD-OIA.
Methods FR $\uparrow$ COUT $\uparrow$ LPIPS $\downarrow$ S3 $\uparrow$ FID $\downarrow$ sFID $\downarrow$
DiME 60.50% -0.4865 0.1119 0.9638 62.3 225.17
ACE 99.90% 0.7278 0.0318 0.9956 9.5 91.70
Ours 99.50% 0.7745 0.0449 0.9949 15.3 92.35
@@
We then present the qualitative analysis. Since ACE achieves the strongest performance among all baselines, we focus our comparison on the counterfactual images generated by our method and those produced by ACE. As shown in Figure 5, similar to the results on the CelebA dataset, the counterfactual images generated by ACE contain many unintended modifications. These include changes in teeth and chin color, the appearance of exaggerated facial creases, and the emergence of extraneous patches on the cheeks. In the transition from “old” to “young”, the counterfactual images generated by our method exhibit comparable visual quality to those produced by ACE, with no obvious undesired changes in either case. As a result, we do not highlight any specific regions for this comparison. However, our method aligns more closely with the intended transformation—for example, by effectively removing age spots and smoothing wrinkles—demonstrating greater semantic consistency with the target concept. In the transition from ”male” to ”female”, we omit the highlight for the same reason.
The extended qualitative results on CelebA, CelebA-HQ, BDD-OIA, and ImageNet are presented in Figures 5 to Figure 10. On CelebA and CelebA-HQ, we highlight the unintended alterations present in the counterfactual images generated by ACE, while also marking the corresponding regions where our method yields noticeable quality improvements. For BDD-OIA and ImageNet, since the visual differences between our method and the baseline are minor, we only present the counterfactual images generated by our method, with the regions of change relative to the original images highlighted.
@@
Figure 5: Qualitative results on CelebA-HQ. Red circles highlight unintended changes produced by ACE, while green circles indicate the corresponding improvements made by our method.
@@
Figure 6: Extended qualitative results on CelebA. Red circles highlight unintended changes produced by ACE, while green circles indicate the corresponding improvements made by our method.
@@
Figure 7: Qualitative results on BDD-OIA. Green circles highlight the counterfactual changes. The right column for each transition direction shows a zoomed-in version.
@@
Figure 8: Extended qualitative results on BDD-OIA. Green circles highlight the counterfactual changes. The right column for each transition direction shows a zoomed-in version.
Figure 9: Qualitative results on ImageNet. Green circles highlight the counterfactual changes.
Figure 10: Extended qualitative results on ImageNet. Green circles highlight the counterfactual changes.
Appendix C VLM Prompt Design
@@
CelebA Smile You are given two sets of facial images: one set contains the original images, and the other contains the corresponding counterfactual images. The target attribute to be modified is smile. If the person in the original image is smiling, the target is ”not smiling”; if the person is not smiling, the target is ”smiling”. Please evaluate the counterfactual images based on the following four criteria, each rated on a scale from 1 to 10:
@@
CelebA Age You are given two sets of facial images: one set contains the original images, and the other contains the corresponding counterfactual images. The target attribute to be modified is age. If the person in the original image is old, the target is ”young”; if the person is young, the target is ”old”. Please evaluate the counterfactual images based on the following four criteria, each rated on a scale from 1 to 10:
  • Validity: Does the counterfactual image successfully transition to the target attribute (i.e., young or old) in a visually convincing way?
CelebA Gender You are given two sets of facial images: one set contains the original images, and the other contains the corresponding counterfactual images. The target attribute to be modified is age. If the person in the original image is male, the target is ”female”; if the person is female, the target is ”male”. Please evaluate the counterfactual images based on the following four criteria, each rated on a scale from 1 to 10:
  • Validity: Does the counterfactual image successfully transition to the target attribute (i.e., male or female) in a visually convincing way?
CelebA-HQ Smile You are given two sets of facial images: one set contains the original images, and the other contains the corresponding counterfactual images. The target attribute to be modified is smile. If the person in the original image is smiling, the target is ”not smiling”; if the person is not smiling, the target is ”smiling”. Please evaluate the counterfactual images based on the following four criteria, each rated on a scale from 1 to 10:
@@
CelebA-HQ Age You are given two sets of facial images: one set contains the original images, and the other contains the corresponding counterfactual images. The target attribute to be modified is age. If the person in the original image is old, the target is ”young”; if the person is young, the target is ”old”. Please evaluate the counterfactual images based on the following four criteria, each rated on a scale from 1 to 10:
  • Validity: Does the counterfactual image successfully transition to the target attribute (i.e., young or old) in a visually convincing way?
@@
CelebA-HQ Gender You are given two sets of facial images: one set contains the original images, and the other contains the corresponding counterfactual images. The target attribute to be modified is age. If the person in the original image is male, the target is ”female”; if the person is female, the target is ”male”. Please evaluate the counterfactual images based on the following four criteria, each rated on a scale from 1 to 10:
  • Validity: Does the counterfactual image successfully transition to the target attribute (i.e., male or female) in a visually convincing way?
@@
  • Validity: Does the counterfactual image successfully transition to the target attribute (i.e., stop or moving forward) in a visually convincing way?
@@
Sorrel - Zebra You are given two sets of animal images: one set contains the original images, and the other contains the corresponding counterfactual images. The target attribute to be modified is animal type. The query attribute is ”sorrel”, and the target attribute is ”zebra”. Please evaluate the counterfactual images based on the following four criteria, each rated on a scale from 1 to 10:
  • Validity: Does the counterfactual image successfully transition to the target attribute (i.e., zebra) in a visually convincing way?
@@
Cougar - Cheetah You are given two sets of animal images: one set contains the original images, and the other contains the corresponding counterfactual images. The target attribute to be modified is animal type. The query attribute is ”cougar”, and the target attribute is ”cheetah”. Please evaluate the counterfactual images based on the following four criteria, each rated on a scale from 1 to 10:
  • Validity: Does the counterfactual image successfully transition to the target attribute (i.e., cheetah) in a visually convincing way?
@@
Persian Cat - Egyptian Cat You are given two sets of animal images: one set contains the original images, and the other contains the corresponding counterfactual images. The target attribute to be modified is animal type. The query attribute is ”Persian cat”, and the target attribute is ”Egyptian cat”. Please evaluate the counterfactual images based on the following four criteria, each rated on a scale from 1 to 10:
  • Validity: Does the counterfactual image successfully transition to the target attribute (i.e., Egyptian cat) in a visually convincing way?
@@
Appendix D Additional Parameters Study
In this section, we provide the parameter study plots for all datasets other than those in Section 5.5, where the observed patterns are consistent with those discussed in the main text. This highlights the critical role of $\alpha$ in our method, as it controls the degree of causal constraint on the change of spurious factors.
Figure 11: Line plots of FR and COUT with respect to varying values of $\alpha$ on CelebA Age.
Figure 12: Line plots of FR and COUT with respect to varying values of $\alpha$ on CelebA Gender.
Figure 13: Line plots of FR and COUT with respect to varying values of $\alpha$ on CelebA-HQ Smile.
Figure 14: Line plots of FR and COUT with respect to varying values of $\alpha$ on CelebA-HQ Age.
Figure 15: Line plots of FR and COUT with respect to varying values of $\alpha$ on CelebA-HQ Gender.
Figure 16: Line plots of FR and COUT with respect to varying values of $\alpha$ on BDD-OIA.
Figure 17: Line plots of FR and COUT with respect to varying values of $\alpha$ on ImageNet (Egyptian Cat - Persian Cat).
Figure 18: Line plots of FR and COUT with respect to varying values of $\alpha$ on ImageNet (Sorrel - Zebra).
Figure 19: Line plots of FR and COUT with respect to varying values of $\alpha$ on ImageNet (Cougar - Cheetah).